OpenStudy (anonymous):

is the sequence geometric? if so, identify the common ratio. 1/3,2/9,4/27,8/81,16/243,....

OpenStudy (ranga):

Yes, the sequence is geometric. The second term divided by the first term is: \[2/9 \div 1/3 = 2/9 \times 3/1 = 2/3\] The third term divided by the second term is: \[4/27 \div 2/9 = 4/27 \times 9/2 = 2/3\] You can try the fourth term divided by the third term which will also be 2/3. Therefore, the sequence is geometric and the common ratio is 2/3.